OpenStudy (anonymous):

A record of travel along a straight path is as follows: 1. Start from rest with a constant acceleration of 2.77 m/s^2 for 15.0 s. 2. Maintain a constant velocity for the next 2.05 min. 3. Apply a constant negative acceleration of 29.47 m/s^2 for 4.39 s. (a) What was the total displacement for the trip? (b) What were the average speeds for legs 1, 2, and 3 of the trip, as well as for the complete trip?

OpenStudy (shamim):

For first 15sec Given u=0 a=2.77m/s t=15sec S1=? We know S1=ut+(1/2)at^2 S1=?

OpenStudy (shamim):

Response plz!!!

OpenStudy (shamim):

I wanna do 2,3 for u

OpenStudy (anonymous):

1/2at^2?

OpenStudy (shamim):

This is (1/2)*a*t^2=0.5*a*t^2 Here a=acceleration t=time

OpenStudy (shamim):

Anyway dont u know this equation s=ut+(1/2)at^2

OpenStudy (shamim):

Can i start 2 number question?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

S1=ut+(1/2)at^2 i've got 312 m/s. is it correct?

OpenStudy (shamim):

I think u r correct

OpenStudy (shamim):

2. Given t=2.05*60sec Required equation is v=u+at=0+2.77*15=? Now S=vt=2.77*15*2.05*60m=?

OpenStudy (anonymous):

=5110.65? *can't open my account -_-

OpenStudy (shamim):

Did u get the above solution?

OpenStudy (shamim):

3. S=ut-.5at^2 S=2.77*15*4.39-.5*29.47*4.39*4.39 S=?

OpenStudy (shamim):

Now add s of 1,2and3

OpenStudy (shamim):

a) U will get total distance travelled

OpenStudy (shamim):

Response plz